# A Preference-Driven Methodology for High-Quality Solidity Code Generation

### v2

@@ Abstract
Index Terms:
Smart Contract Generation, Supervised Fine-Tuning, Direct Preference Optimization
11footnotetext: Equal contribution.22footnotetext: Corresponding authors.
@@ I Introduction
Equally critical is the gap in gas efficiency optimization, which represents another dimension of human preference that current LLMs largely overlook. Gas efficiency is a paramount concern for Ethereum deployment, where execution costs directly impact user adoption and economic viability [5]. While models like GPT-4 [6] and DeepSeek-R1 [7] excel at producing functionally correct contracts, they typically neglect gas consumption optimization, resulting in contracts that meet functional requirements but violate user preferences for cost-effectiveness [8, 2]. The fundamental issue lies in the misalignment between current optimization objectives and genuine human preferences for smart contracts. Optimizing for one dimension often sacrifices others [8, 9], creating a critical trade-off that is particularly problematic in decentralized finance (DeFi) applications, where high gas costs can undermine economic feasibility [10, 11]. Although recent advances leverage Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) for code preference alignment, existing approaches address correctness, efficiency, and security in isolation [2, 9], failing to capture the holistic preference optimization required for production-ready smart contracts that satisfy real-world user requirements [8].
@@
  • Production-Ready Smart Contract Generation: PrefGen achieves breakthrough performance improvements: 66.7% Pass@5, 58.9% Gas@5, and 62.5% Secure@5 on Qwen-7B, with practical deployment benefits including 12% reduced gas costs in ERC-20 contracts and complete elimination of critical security vulnerabilities such as reentrancy attacks.
Figure 1: Overview of the PrefGen framework v.s SFT + DPO.
@@ II-A3 SFT and DPO
Recent work has demonstrated the effectiveness of combining SFT and DPO for code generation tasks [26], showing that this two-phase approach enables models to first establish functional competence through SFT, then refine quality through preference optimization. For smart contract generation, this progression is particularly critical: establishing basic Solidity syntax and semantics through SFT creates a foundation for subsequent optimization of blockchain-specific objectives through DPO.
@@ III Approach
This section presents PrefGen, a framework leverage multi-objective optimization to generate functionally correct, gas-efficient, and secure Solidity smart contracts.
@@ III-B Data Quality Ranking Algorithm
where $d$ is the damping factor and $\text{Link}(t,c)$ indicates whether code snippet $c$ passes test case $t$. This iterative refinement converges to rankings that accurately reflect code quality based on correctness validation. Higher-ranked implementations are prioritized during training, ensuring the model learns from verified, high-quality exemplars.
@@ III-D2 Multi-Objective Optimization
where $\alpha$ and $\beta$ are weights of the corresponding rewards.
@@ IV-A Constructed Dataset
Stage 1: Base Dataset Collection. We utilize SolEval [2] as our foundation and expand it by randomly crawling 19 high-star GitHub projects, resulting in 1,125 function samples for training and 382 samples for testing. This expansion ensures broader coverage of real-world smart contract patterns and implementation diversity.
Stage 2: Multi-Model Code Generation. Following SolEval’s best practices and the prompt construction pipeline shown in Figure 2, we perform inference across 16 mainstream large language models to generate 104,640 function samples. We then apply rigorous filtering to remove functionally incorrect implementations, retaining 12,096 valid function samples that pass all unit tests.
Stage 3: Preference-Based Sample Selection. We employ PageRank algorithm to rank and select high-quality samples for preference optimization. This process yields our final multi-objective dataset comprising 6,520 functional-correctness optimization samples, 533 security-related samples, and 533 gas optimization samples, each containing unique problem prompts with preferred and rejected solution pairs.
@@ IV-B Studied Large Language Models
We select 16 state-of-the-art LLMs widely used in recent code generation studies [31, 32, 33, 34, 35]. In particular, we focus on recent models released since 2022, and we exclude the small models (with fewer than 2B parameters) due to their limited efficacy. Table II presents the state-of-the-art LLMs studied in our experiments with their sizes and types. For RQ-2 and RQ-3, we choose Qwen-7B due to its strong performance in Table III, making it a promising candidate for further fine-tuning. Parameter count constraints are imposed by our available computational resources (8 × NVIDIA A800 GPUs).
TABLE II: Overview of the studied LLMs
@@ IV-C Evaluation Metrics
[…]
@@ IV-D Implementation
Training Hyperparameters. Our PrefGen framework employs a multi-stage training approach. For Supervised Fine-Tuning (SFT), we train for 3 epochs with a maximum input length of 2048 tokens, using a 9:1 training/validation split. For the multi-objective optimization phase, we set hyperparameters $\alpha=\beta=1.0$ for balanced security and efficiency optimization, and $\lambda=0.5$ for the security component based on empirical validation. All fine-tuning experiments use a batch size of 2 and learning rate of 1e-5.
Evaluation Setup. For all experiments, we generate $n=10$ samples per problem and evaluate using the unbiased estimator for Task@k metrics with $k\leq 10$. Temperature is set to 0.8 for diverse code generation while maintaining quality. Gas efficiency is measured using Forge execution, and security analysis employs Slither [18] for vulnerability detection.
V Results
@@ V-B RQ-2 Framework Effectiveness Evaluation
Effectiveness Analysis. As shown in Table IV, PrefGen demonstrates superior performance across all evaluation metrics compared to traditional fine-tuning approaches. PrefGen achieves the highest Pass@5 (66.7%), significantly outperforming SFT (58.3%), DPO (25.0%), and SFT+DPO (55.7%). More notably, PrefGen excels in the challenging multi-objective optimization scenario, achieving Gas@5 of 58.9% and Secure@5 of 62.5%, substantially surpassing all baseline methods. The results demonstrate that PrefGen’s multi-objective training approach effectively balances correctness, efficiency, and security, while traditional sequential approaches (SFT + DPO) show trade-offs between these objectives.
@@
Efficiency Analysis. Table V reveals that PrefGen achieves superior training efficiency despite higher memory overhead. While PrefGen requires 193.11 GB memory compared to SFT’s 76.84 GB, it completes training in only 8:53, dramatically outperforming SFT (1:22:35) and SFT+DPO (1:33:46) by 90% reduction in training time. The substantial time savings, coupled with comprehensive performance improvements across correctness, gas efficiency, and security, demonstrate PrefGen’s practical viability for production smart contract development.
@@
Answer to RQ-2: PrefGen significantly outperforms all baseline approaches (SFT, DPO, SFT+DPO) across all metrics, achieving 66.7% Pass@5, 58.9% Gas@5, and 62.5% Secure@5. While PrefGen has moderately higher computational cost than SFT alone (approximately equivalent to DPO, lower than SFT+DPO), this efficiency trade-off is well-justified by the substantial improvements in correctness, gas efficiency, and security. PrefGen effectively balances all three critical dimensions, making PrefGen highly suitable for real-world smart contract deployment.
@@ VI Threats to Validity
Internal Validity. Our dataset derives from existing Solidity repositories, potentially introducing bias toward specific coding patterns. To mitigate this threat, we implement a three-stage dataset construction process (Table I) that systematically expands beyond a single source: (1) we crawl 19 diverse high-star GitHub projects spanning different domains (security, economics, games), (2) we generate 104,640 samples using 16 different LLMs to ensure implementation diversity, and (3) we employ PageRank algorithm for preference-based sample selection, yielding 7,586 balanced preference pairs across correctness, security, and gas optimization. Gas measurements utilize Ethereum’s current pricing model through Forge execution, while security analysis employs Slither [18].
@@
Construct Validity. Our metrics may not capture all aspects of smart contract quality, potentially missing nuanced developer preferences. To strengthen construct validity, we implement a unified Task@k evaluation framework that extends beyond traditional Pass@k to include Compile@k, Gas@k, and Secure@k, providing comprehensive multi-dimensional assessment. Our RQ-2 framework evaluation (Table IV) demonstrates that PrefGen achieves balanced improvements across all three dimensions (66.7% Pass@5, 58.9% Gas@5, 62.5% Secure@5), validating our multi-objective approach. We use unbiased estimators from HumanEval [36] to ensure statistical rigor ($n=10$ samples per problem). Real-world case studies with ERC-20 and ERC-721 contracts (Listing 1) provide qualitative validation of practical improvements.
@@ VII-A Code Generation Framework
The advancement of pre-training technology has significantly advanced code generation in both academia and industry [40, 41, 42, 43]. This has led to the emergence of numerous Large Language Models (LLMs) that have made substantial strides in code generation [44, 45, 46, 47, 48, 49]. Recent advances have introduced more sophisticated models [50, 7, 51, 52]. These models showcase the evolution from simple token prediction to understanding complex code structures and dependencies.
To optimize LLMs for various code generation scenarios, previous studies have focused on enhancing prompt engineering by introducing specific patterns, such as Structured Chain-of-Thought [53, 54], Self-planning [55], Self-debug [56, 57], and Self-collaboration [58, 59]. However, these efforts primarily address mainstream programming languages (e.g., Java, Python, and C++) [59, 60, 57], with limited attention to domain-specific languages like Solidity. Repository-level code generation has emerged as a critical area of research. A3-CodGen [33] proposes a framework that considers local, global, and third-party library awareness for code reuse. Similarly, Shrivastava et al. [61] introduce repository-level prompt generation techniques for LLMs. These approaches underscore the critical role of contextual understanding in smart contract generation, where inter-contract interactions are prevalent. Beyond traditional benchmarks, recent evaluation frameworks have emerged. LiveCodeBench [62] addresses contamination issues in code generation evaluation by using continuously updated problems. CoderEval [34] focuses on pragmatic code generation scenarios, while Evocodebench [35] evaluates code generation in practical software projects. ClassEval [63] specifically targets class-level code generation, moving beyond simple function generation.
@@ VII-B Defect Detection and Gas Fee Optimization
Recent work has begun leveraging LLMs for smart contract analysis. Jiang et al. [20] demonstrate the potential of LLMs in identifying gas-wasting code smells in smart contracts, achieving promising results in detecting inefficient patterns that lead to unnecessary gas consumption. This approach combines the pattern recognition capabilities of LLMs with domain-specific knowledge about gas optimization. LLM-based approaches have shown promise in understanding complex code patterns. Studies have explored how LLMs handle ambiguous code contexts [66] and how they can be distracted by irrelevant information [67]. Jin et al. [68] propose methods to extend LLM context windows without fine-tuning, which is particularly relevant for analyzing large smart contract codebases. These insights inform the design of more effective smart contract analysis tools. Gas optimization remains critical in smart contract development, with established strategies including storage packing, loop optimization, and redundant operation elimination [27, 69, 70]. However, existing tools [69, 70, 71] primarily focus on post-development optimization rather than generating inherently gas-efficient code. This reactive approach creates a fundamental gap that motivates the need for code generation frameworks with built-in gas efficiency considerations. Integrating defect detection and gas optimization into the code generation pipeline represents a paradigm shift from traditional sequential approaches [72, 73]. While conventional methods treat security analysis and efficiency optimization as separate post-generation phases [72, 73], incorporating these considerations directly into the generation process enables the production of inherently robust and cost-effective smart contracts. This proactive integration is crucial given the immutable nature of blockchain deployments [12, 13, 14, 21, 22], where vulnerabilities cannot be patched post-deployment and have led to billions of dollars in financial losses [15, 16, 17].

### v3

# PrefGen: A Preference-Driven Methodology for Secure Yet Gas-Efficient Smart Contract Generation

@@ Abstract
22footnotetext: Equal contribution.11footnotetext: Corresponding authors.
Chao Ni is also with Hangzhou High-Tech Zone (Binjiang) Institute of Blockchain and Data Security.
@@ I Introduction
Equally critical is the gap in gas efficiency optimization, which represents another dimension of human preference that current LLMs largely overlook. Gas efficiency is a paramount concern for Ethereum deployment, where execution costs directly impact user adoption and economic viability [5, 6]. While models like GPT-4 [7] and DeepSeek-R1 [8] excel at producing functionally correct contracts, they typically neglect gas consumption optimization, resulting in contracts that meet functional requirements but violate user preferences for cost-effectiveness [9, 2]. The fundamental issue lies in the misalignment between current optimization objectives and genuine human preferences for smart contracts. Optimizing for one dimension often sacrifices others [9, 10], creating a critical trade-off that is particularly problematic in decentralized finance (DeFi) applications, where high gas costs can undermine economic feasibility [11, 12]. Although recent advances leverage Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) for code preference alignment, existing approaches address correctness, efficiency, and security in isolation [2, 10], failing to capture the holistic preference optimization required for production-ready smart contracts that satisfy real-world user requirements [9].
@@
  • Production-Ready Smart Contract Generation: PrefGen achieves breakthrough performance improvements: 66.7% Pass@5, 58.9% Gas@5, and 62.5% Secure@5 on Qwen-7B, with practical deployment benefits including 12% reduced gas costs in ERC-20 contracts and complete elimination of critical security vulnerabilities such as reentrancy attacks. We open-source our dataset and evaluation framework at [13].
@@ II-A3 SFT and DPO
Recent work has demonstrated the effectiveness of combining SFT and DPO for code generation tasks [28, 31], showing that this two-phase approach enables models to first establish functional competence through SFT, then refine quality through preference optimization. For smart contract generation, this progression is particularly critical: establishing basic Solidity syntax and semantics through SFT creates a foundation for subsequent optimization of blockchain-specific objectives through DPO.
@@ III Approach
This section presents PrefGen, a framework that leverages multi-objective optimization to generate functionally correct, gas-efficient, and secure Solidity smart contracts.
@@ III-B Data Quality Ranking Algorithm
where $d$ is the damping factor and $\text{Link}(t,c)$ indicates whether code snippet $c$ passes test case $t$. Following established PageRank practices [32, 28], we use $d=0.85$. Our ablation analysis demonstrates this choice: at $d=0.50$, the model achieves 63.1% Pass@5, 52.4% Gas@5, and 56.0% Secure@5; increasing to $d=0.65$ yields 64.8%, 54.3%, 58.2%; at $d=0.75$ produces 66.0%, 58.0%, 61.3%; the canonical $d=0.85$ peaks at 66.7%, 58.9%, 62.5%; while $d=0.95$ scores decline to 65.4%, 57.7%, 60.8%. This iterative refinement converges to rankings that accurately reflect code quality based on correctness validation. Higher-ranked implementations are prioritized during training, ensuring the model learns from verified, high-quality exemplars.
@@ III-D2 Multi-Objective Optimization
where $\alpha$ and $\beta$ are weights of the corresponding rewards. We conduct sensitivity analysis to determine optimal values: at $\alpha=\beta=1$, the model achieves 66.7% Pass@5, 58.9% Gas@5, and 62.5% Secure@5. For security-first settings ($\alpha=1.2,\beta=0.8$), performance shifts to 64.3% Pass@5, 52.1% Gas@5, and 69.8% Secure@5. Conversely, with cost-first settings ($\alpha=0.8,\beta=1.2$), the method yields 65.9% Pass@5, 64.2% Gas@5, and 58.7% Secure@5, demonstrating effective multi-objective trade-offs.
@@ IV-A Constructed Dataset
Stage 1: Base Dataset Collection. We utilize SolEval [2] as our foundation and expand it by systematically selecting 19 high-star GitHub projects based on specific criteria: star ratings, recent commits (within 6 months), diverse contract types (e.g., DeFi, NFT), and test case availability. We specifically filter out focal methods lacking test cases, resulting in an average of 2.4 test cases per focal function. This systematic selection yields 1,125 function samples for training and 382 samples for testing, ensuring broader coverage of real-world smart contract patterns and implementation diversity.
Stage 2: Multi-Model Code Generation. Following SolEval’s best practices and the prompt construction pipeline shown in Figure 2, we perform inference across 16 mainstream large language models to generate 104,640 function samples. To handle repository-level context, we first retrieve source files and parse them with Treesitter to extract types, functions, state variables, and constants. We employ static program analysis to identify contextual dependencies by capturing all function calls defined outside the current function scope, extracting their signatures, and building a dependency graph connecting the focal function to its required context. We store the invocation signatures along with their definitions, creating a comprehensive context database for each function sample. We then apply rigorous filtering to remove functionally incorrect implementations, retaining 12,096 valid function samples that pass all unit tests.
Stage 3: Preference-Based Sample Selection. We employ the PageRank algorithm to rank and select high-quality samples for preference optimization. This process yields our final multi-objective dataset comprising 6,520 functional-correctness optimization samples, 533 security-related samples, and 533 gas optimization samples, each containing unique problem prompts with preferred and rejected solution pairs.
@@ IV-B Studied Large Language Models
We select 16 state-of-the-art LLMs widely used in recent code generation studies [34, 35, 36, 37, 38]. In particular, we focus on recent models released since 2022, and we exclude the small models (with fewer than 2B parameters) due to their limited efficacy. Table II presents the state-of-the-art LLMs studied in our experiments with their sizes and types. Our experiments are based on Qwen2.5-Coder, released on November 12, 2024, ensuring evaluation with current state-of-the-art capabilities. For RQ-2 and RQ-3, we choose Qwen2.5-Coder-7B due to its strong performance in Table III, making it a promising candidate for further fine-tuning. Parameter count constraints are imposed by our available computational resources (8 × NVIDIA A800 GPUs).
TABLE II: Overview of the studied LLMs.
@@ IV-C Evaluation Metrics
$\text{Task@}k:=\mathop{\mathbb{E}}\limits_{\text{Problems}}\left[1-\frac{\binom{n-c_{\text{task}}}{k}}{\binom{n}{k}}\right$] (8)
@@ IV-D Implementation
Training Hyperparameters. Our PrefGen framework employs a multi-stage training approach. For Supervised Fine-Tuning (SFT), we train for 3 epochs with a maximum input length of 2048 tokens, using a 9:1 training/validation split. For the multi-objective optimization phase, we set hyperparameters $\alpha=\beta=1.0$ for balanced security and efficiency optimization, and $\lambda=0.5$ for the security component based on empirical validation. All fine-tuning experiments use a batch size of 2 and a learning rate of 1e-5.
Evaluation Setup. For all experiments, we generate $n=10$ samples per problem and evaluate using the unbiased estimator for Task@k metrics with $k\leq 10$. Temperature is set to 0.8 for diverse code generation while maintaining quality. For Pass@k evaluation, we employ Forge to conduct comprehensive fuzzing on test cases within a simulated blockchain environment, where functions failing to pass fuzzing are counted as failures in Pass@k calculation. Gas efficiency is measured using Forge execution, and security analysis employs Slither [20] for vulnerability detection.
@@
Performance Metrics. Inference latency analysis reveals: Qwen-2.5-Coder-7B baseline requires 2.74s per sample, SFT achieves 1.46s per sample, SFT+DPO requires 1.79s per sample, while PrefGen requires 1.98s per sample. Training costs are estimated using AWS deployment: fully loaded on-demand training of Qwen-7B with PrefGen on AWS costs approximately \$7.6 per epoch using p4d.24xlarge instances with 8×A100-40GB GPUs.
V Results
@@ V-A RQ-1 Empirical Evaluation
To validate PrefGen’s broader applicability, we further evaluate newer models. Qwen3-8B shows limited zero-shot performance with 7.2% Pass@5, 1.7% Gas@5, and 4.3% Secure@5, which significantly improves to 28.9% Pass@5, 9.3% Gas@5, and 5.8% Secure@5 after SFT. With SFT+DPO, it achieves 26.9% Pass@5, 17.6% Gas@5, and 11.3% Secure@5, while PrefGen further enhances performance to 29.2% Pass@5, 18.9% Gas@5, and 12.1% Secure@5. Similarly, DeepSeek-R1-Distill-Qwen-7B demonstrates substantial improvements: from 4.5% Pass@5, 5.3% Gas@5, and 1.0% Secure@5 in zero-shot to 22.1% Pass@5, 7.8% Gas@5, and 3.7% Secure@5 with SFT. SFT+DPO achieves 23.8% Pass@5, 9.1% Gas@5, and 8.5% Secure@5, while PrefGen reaches 26.8% Pass@5, 14.2% Gas@5, and 12.5% Secure@5. These results consistently demonstrate PrefGen’s effectiveness across diverse model architectures.
@@ V-B RQ-2 Framework Effectiveness Evaluation
Effectiveness Analysis. As shown in Table IV, PrefGen demonstrates superior performance across all evaluation metrics compared to traditional fine-tuning approaches. PrefGen achieves the highest Pass@5 (66.7%), significantly outperforming SFT (58.3%), DPO (25.0%), and SFT+DPO (55.7%). More notably, PrefGen excels in the challenging multi-objective optimization scenario, achieving Gas@5 of 58.9% and Secure@5 of 62.5%, substantially surpassing all baseline methods. To illustrate PrefGen’s practical advantages, we conducted a targeted experiment with 100 samples comparing post-hoc filtering versus integrated training. Plain SFT produces code that passes all tests in 70 cases but satisfies both gas budget and security checks in only 8 cases. Simple post-hoc filtering (20% threshold) creates a fundamental trade-off: keeping either the most gas-efficient or the most secure outputs results in only 2 contracts remaining fully compliant across all dimensions. In contrast, PrefGen trains the model to inherently balance objectives rather than trimming outputs post-generation. After two multi-objective DPO epochs, 82 generations pass functionality tests while 65 simultaneously meet both gas and security requirements—an eight-fold improvement in deployable yield. This demonstrates that PrefGen’s PageRank-guided positive/negative sampling enables models to learn multi-objective preferences simultaneously rather than optimizing single objectives in isolation.
@@
Efficiency Analysis. Table V reveals that PrefGen achieves superior training efficiency despite higher memory overhead. While PrefGen requires 193.11 GB of memory compared to SFT’s 76.84 GB, it completes training in only 8:53, dramatically outperforming SFT (1:22:35) and SFT+DPO (1:33:46) by 90% reduction in training time. The substantial time savings, coupled with comprehensive performance improvements across correctness, gas efficiency, and security, demonstrate PrefGen’s practical viability for production smart contract development.
@@
Answer to RQ-2: PrefGen significantly outperforms all baseline approaches (SFT, DPO, SFT+DPO) across all metrics, achieving 66.7% Pass@5, 58.9% Gas@5, and 62.5% Secure@5. While PrefGen has a moderately higher computational cost than SFT alone (approximately equivalent to DPO, lower than SFT+DPO), this efficiency trade-off is well-justified by the substantial improvements in correctness, gas efficiency, and security. PrefGen effectively balances all three critical dimensions, making PrefGen highly suitable for real-world smart contract deployment.
@@ VI Threats to Validity
Internal Validity. Our dataset derives from existing Solidity repositories, potentially introducing bias toward specific coding patterns. To mitigate this threat, we implement a three-stage dataset construction process (Table I) that systematically expands beyond a single source: (1) We crawl 19 diverse high-star GitHub projects spanning different domains (security, economics, games), (2) We generate 104,640 samples using 16 different LLMs to ensure implementation diversity, and (3) We employ the PageRank algorithm for preference-based sample selection, yielding 7,586 balanced preference pairs across correctness, security, and gas optimization. Gas measurements utilize Ethereum’s current pricing model through Forge execution, while security analysis employs Slither [20].
@@
Construct Validity. Our metrics may not capture all aspects of smart contract quality, potentially missing nuanced developer preferences. To strengthen construct validity, we implement a unified Task@k evaluation framework that extends beyond traditional Pass@k to include Compile@k, Gas@k, and Secure@k, providing a comprehensive multi-dimensional assessment. Our RQ-2 framework evaluation (Table IV) demonstrates that PrefGen achieves balanced improvements across all three dimensions (66.7% Pass@5, 58.9% Gas@5, 62.5% Secure@5), validating our multi-objective approach. We use unbiased estimators from HumanEval [39] to ensure statistical rigor ($n=10$ samples per problem). Real-world case studies with ERC-20 and ERC-721 contracts (Listing 1) provide qualitative validation of practical improvements.
@@ VII-A Code Generation Framework
The advancement of pre-training technology has significantly advanced code generation in both academia and industry [43, 44, 45]. This has led to the emergence of numerous Large Language Models (LLMs) that have made substantial strides in code generation [46, 47, 48, 49, 50, 51]. Recent advances have introduced more sophisticated models [52, 8, 53]. These models showcase the evolution from simple token prediction to understanding complex code structures and dependencies.
To optimize LLMs for various code generation scenarios, previous studies have focused on enhancing prompt engineering by introducing specific patterns, such as Structured Chain-of-Thought [54, 55], Self-planning [56], Self-debug [57, 58], and Self-collaboration [59]. However, these efforts primarily address mainstream programming languages (e.g., Java, Python, and C++) [60, 61, 58], with limited attention to domain-specific languages like Solidity. Repository-level code generation has emerged as a critical area of research. A3-CodGen [36] proposes a framework that considers local, global, and third-party library awareness for code reuse. Similarly, Shrivastava et al. [62] introduce repository-level prompt generation techniques for LLMs. These approaches underscore the critical role of contextual understanding in smart contract generation, where inter-contract interactions are prevalent. Beyond traditional benchmarks, recent evaluation frameworks have emerged. LiveCodeBench [63] addresses contamination issues in code generation evaluation by using continuously updated problems. CoderEval [37] focuses on pragmatic code generation scenarios, while Evocodebench [38] evaluates code generation in practical software projects. ClassEval [64] specifically targets class-level code generation, moving beyond simple function generation.
@@ VII-B Defect Detection and Gas Fee Optimization
Recent work has begun leveraging LLMs for smart contract analysis. Jiang et al. [22] demonstrate the potential of LLMs in identifying gas-wasting code smells in smart contracts, achieving promising results in detecting inefficient patterns that lead to unnecessary gas consumption. This approach combines the pattern recognition capabilities of LLMs with domain-specific knowledge about gas optimization.
Existing preference learning approaches for code generation, including CodeDPO [28] and SafeDPO [67], primarily focus on single objectives among functional correctness, gas efficiency, and security, lacking unified multi-objective optimization frameworks. CodeDPO and Focused-DPO improve general code correctness through preference learning but do not address the unique challenges of smart contract generation. Our work introduces a fundamental paradigm shift by incorporating quantifiable blockchain-specific rewards that enable simultaneous optimization of all three critical dimensions—functional correctness, gas efficiency, and security—within a unified preference learning framework specifically designed for smart contract generation. LLM-based approaches have shown promise in understanding complex code patterns. Studies have explored how LLMs handle ambiguous code contexts [68] and how they can be distracted by irrelevant information [69]. Jin et al. [70] propose methods to extend LLM context windows without fine-tuning, which is particularly relevant for analyzing large smart contract codebases. These insights inform the design of more effective smart contract analysis tools. Gas optimization remains critical in smart contract development, with established strategies including storage packing, loop optimization, and redundant operation elimination [29, 71, 72]. However, existing tools [71, 72] primarily focus on post-development optimization rather than generating inherently gas-efficient code. This reactive approach creates a fundamental gap that motivates the need for code generation frameworks with built-in gas efficiency considerations. Integrating defect detection and gas optimization into the code generation pipeline represents a paradigm shift from traditional sequential approaches [73, 74]. While conventional methods treat security analysis and efficiency optimization as separate post-generation phases [73, 74], incorporating these considerations directly into the generation process enables the production of inherently robust and cost-effective smart contracts. This proactive integration is crucial given the immutable nature of blockchain deployments [14, 15, 16, 23, 24], where vulnerabilities cannot be patched post-deployment and have led to billions of dollars in financial losses [17, 18, 19].